OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a14 = -33, a15 = 9 an = -579 + 42(n + 1) an = -579 + 42(n - 1) an = -579 - 42(n + 1) an = -579 - 42(n - 1)

OpenStudy (chaise):

Just plug in the value n=14 and n=15 in to all of those equations, and the if you get -33 and 9 respectively, that is the answer that is correct. Do you need help with how to generate an arithmatic sequence based upon two consecutive terms?

OpenStudy (anonymous):

i need help with this whole lesson to be honest.. ive been really struggling

OpenStudy (welshfella):

nth term = a1 + (n - 1)d where a1 = first term, and d = common difference in this case the value of d = 9 - (-33)

OpenStudy (welshfella):

you can find what the first term is by substituting in the formula an = a1 +(n-1)d

OpenStudy (welshfella):

what is the value of d 9 -(-33) = 42 right?

OpenStudy (anonymous):

yes

OpenStudy (welshfella):

15th term = 9 so substituting 9 = a1 + (15 - 1)42

OpenStudy (welshfella):

from this we can find the first term a1

OpenStudy (welshfella):

a1 = 9 - 14*42 = -579

OpenStudy (welshfella):

get it or have i gone too fast?

OpenStudy (anonymous):

okay that sounds easy.....

OpenStudy (welshfella):

yes you just plug in the values for a1 and d into an = a1 + d(n - 1)

OpenStudy (welshfella):

Arithmetic sequences are of the form a1 , a1+d, a1 + 2d, a1 + 3d ....... each value after the first is obtained by adding the common difference d

OpenStudy (welshfella):

a simple one would be 3 , 6, 9, 12 ....

OpenStudy (anonymous):

im so sorry.. its really just not clicking together right now.. okay so for that example you just gave ... 3= a1 correct? then a1+d= 6? and so on .. right??

OpenStudy (welshfella):

yes second term is 3 +3 = 6

OpenStudy (welshfella):

then 3 + 2(3) = 9

OpenStudy (welshfella):

glad to be of help but i just got to go right now

OpenStudy (anonymous):

thanks for the help ! :)

OpenStudy (anonymous):

SO the answer would be B an = -579 + 42(n - 1)